OpenStudy (anonymous):

If y = 3x^3 – 2x and dx/dt = 3, find dy/dt when x = –2.

OpenStudy (anonymous):

@hartnn @Luigi0210 @paki @Destinymasha @confluxepic @sammixboo @Nnesha @undeadknight26 @eliassaab

OpenStudy (anonymous):

\[\frac {dy}{dt} = 6 x^2 \frac {dx}{dt} - 2 \frac {dx}{dt} \]

OpenStudy (anonymous):

You have x= -2 and dx/dt=3, replace and you will be done.

OpenStudy (anonymous):

102?

OpenStudy (anonymous):

its 9x^2-2

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

\[ 6 (-2)^2 (-3)- 2(-3)=-66 \]